OpenStudy (anonymous):

anyone ever seen this: integral of 2x+3/sqrt(4x-x^2) what is the substitution? Do you seperate into 2 parts?

OpenStudy (anonymous):

When no one is around that knows calculus I go to www.WolframAlpha.com http://www.wolframalpha.com/input/?i=integrate+%282x%2B3%29%2Fsqrt%284x-x^2%29+

OpenStudy (mr.math):

Hint: \[{2x+3 \over \sqrt{4x-x^2}}=-{-2x+4 \over \sqrt{4x-x^2}}+{7 \over \sqrt{4x-x^2}}\]

OpenStudy (jamesj):

exactly. And for the second term you'll need to complete the square. WolframAlpha is the double-edged sword for calculus learners. I'd stay away from it and try and figure out problems yourself; you'll learn much, much more that way.

OpenStudy (anonymous):

thank you- both of you i'll work along that course

jhonyy9 (jhonyy9):

- so the first will be x not can be equal zero never ,because than the fraction will be undefined

jhonyy9 (jhonyy9):

and x not can be never equal 4